# ICGNet: A Unified Approach for Instance-Centric Grasping

### v1

@@ I Introduction
In experiments, significant improvements arise from our object-centric architecture, establishing a new state-of-the-art in the packed decluttering benchmark introduced in [4] while surpassing scene-centric task baselines [8, 6]. These improvements are reflected in real-world experiments deployed on the Franka Research 3 robot. We additionally illustrate how instance-centric information can facilitate target-driven grasping (i.e., “grasp instance number 1" or “grasp the bottle") and effectively prevent post-grasp object-object collisions
@@ II Related Work
Deep Grasp Synthesis.   Recent advancements in robotic grasping have predominantly utilized deep learning techniques to detect robotic grasps directly from sensor data. Unlike more traditional heuristics [16], these methods have demonstrated superior generalization performance towards previously unseen objects as well as achieving successful grasps in cluttered scenes. While several methods have achieved high success rates in 4-degree-of-freedom (4 DoF) top-down grasping [2, 17], their capabilities are often limited when dealing with cluttered scenes or task-dependent grasping scenarios [9]. This limitation has led to an increased research focus on 6 DoF grasp predictions, which aim to predict the full grasp pose from visual observations. We further classify these methods in sampling- and scene-based approaches. Sampling-based methods sample different grasp candidates and use a detector network to refine them further. GPD [3] and PointNetGPD [18] learn to detect grasp poses in cluttered scenes from raw pointclouds by first sampling feasible grasp predictions and then using a CNN or PointNet classifier to predict the quality of each grasp based on the enclosed points. 6-DoF Graspnet [9] extends the way that grasp proposals are generated to the full SE(3) space by leveraging a variational auto-encoder for singulated objects. Scene-based approaches directly predict feasible grasp poses for the whole scene in one forward pass. Contact based methods such as Edge Grasp Networks [6] or Contact GraspNet [5] directly predict the SE(3) pose, width, and grasp quality for each point in the pointcloud. VGN [4] and GIGA [8] predict grasps for each voxel in the reconstructed TSDF using 3D CNN.
@@ III Method
Figure 2: Grasp Representation. Left: $\mathbf{c}$ is the contact point of the closed gripper and $\mathbf{n}$ is its estimated surface normal. $\mathbf{a}$ is the approach direction of the gripper. Given the gravity vector $\mathbf{z}$ and surface normal $\mathbf{n}$, $\mathbf{a}$ can be uniquely defined by the approach angle $\alpha$. Top Right: Grasp ambiguity of different gerasp representations. When dealing with a particular contact or gripper center, there can be multiple feasible approach direction resulting in a successful grasp. Bottom Right: For each contact point, our representation enables the prediction of grasp qualities for different gripper orientations perpendicular to the surface normal.
@@
The corresponding set of SE(3) poses for a given grasp affordance prediction $g$ and gravity vector $\mathbf{z}$ is computed as
@@
Encoder. Given a pointcloud in the world frame, our encoder network performs several key tasks. It extracts both sparse and dense features at multiple resolutions utilizing a sparse 3D-UNet architecture [24] for surface features and dense 3D-UNet architecture [25] for volumetric features. The sparse features are enriched with volumetric information through the proposed token aggregation module. To decompose the voxelized scene into individual instances, we apply multiple query refinement modules similar to Mask3D [26]. Further, we rely on a classification head to assign a class to each latent representation and filter out unmatched queries using a no-object class. All valid instance queries are directly used as latent embeddings for an occupancy network. Additionally, we employ a final Self Attention and MLP layer to exchange inter-object information between the queries and convert them to the affordance domain.
@@
Decoder. Our decoder is designed as an implicit neural field and uses world coordinates ($x,y,z$) as input. It predicts occupancy $p_{occ}$ and grasp affordances $g\in\mathcal{G}_{aff}$ for each instance. In addition, predictions such as classification and instance heatmaps are directly extracted from the encoder as shown in Fig. 3. These predictions are made for each query (class) and each point within the input pointcloud (instance id). To model each implicit decoder, we use a series of MLPs with residual connections following the approach proposed in [27]. Further, we concatenate each query with its positional encoding, which captures spatial information. Moreover, the queried coordinates ($x,y,z$) are enriched by incorporating the surface and volumetric features using our interpolation module. These enriched coordinates are then concatenated with the positional encodings of the original coordinates.
@@
Token Aggregation. Ideally, our goal is to extract dense features at a high resolution and directly feed them into the query refinement module. However, this approach comes with significant drawbacks, notably increased memory consumption and computation cost, rendering it infeasible for most applications. On the other hand, sparse neural networks scale with the number of occupied voxels, making them suitable for scaling to large areas. This advantage arises as the number of occupied voxels typically grows slower than the total number of voxels. Here, we combine the best of both by enriching each occupied surface feature with a volumetric context that is extracted from the volumetric feature grid. To this end, each surface feature is concatenated with the feature of the nearest volumetric voxel, similar to PointNet [28]. This enables the volumetric feature backbone to operate on a larger resolution extracting context information to enrich the sparse features used in the query refinement.
@@
Query Refinement. Given a set of $K$ instance queries, we apply a series of masked cross-attention and self-attention to extract instance-centric information given the scene-level features extracted from the U-Net backbones. We adapt Mask3D [26] and add an MLP layer to further process the extracted scene features. We also add Fourier positional encodings [29] based on voxel positions and use farthest-point sampling to sample query positions.
Classifier. The classifier predicts a class label including a non-object class to address the varying number of instances in a scene. A small MLP followed by a softmax activation predicts a categorical distribution over the desired $C+1$ classes. We manually annotate the points with class labels that correlate with the respective shape, consisting of six categories (“mug", “box", “can", “bottle", “cylindric", “ball" and “other"). Having the notion of semantic classes, allows to predict grasp candidates and reconstructions for individual objects as shown in Fig. 4.
@@
Loss Formulation.   The proposed model is trained in an end-to-end fashion. The training uses instance and semantic annotations as well as occupancy and grasp poses obtained from simulation. The implemented loss is given as
@@
where the first three terms refer to the panoptic segmentation task and are calculated with respect to the input pointcloud $P_{\mathtt{scan}}$. $\mathcal{L}^{Grasp}$ refers to the grasp predictions for each instance. This loss is calculated with respect to the sampled grasp coordinates which may differ from the actual observed pointcloud. We use the cross entropy loss to supervise the approach classification and the squared L2 loss for the gripper width. Finally, the reconstruction loss $\mathcal{L}_{\mathtt{occ}}^{BCE}$ supervises the predicted occupancy for each instance and is computed with respect to an additional set of randomly sampled points within the scene. Note that there is no ordering in the set of instances in a scene and we need to establish correspondence between the set of predicted and set of groundtruth instances during training. Following [26], we use Hungarian matching based on the instance segmentation  [^3] We limit the matching to the segmentation task since calculating every possible assignment of the occupancy field is computationally extensive. loss to find unique assignments and apply the same loss function to the predictions at multiple resolutions.
@@
${}^{\dagger}$ [8] 89.9 $\pm$ 1.7 87.6 $\pm$ 2.0 76.3 $\pm$ 2.4 80.9 $\pm$ 4.1 2.5 $\pm$ 1.2 82 $\pm$ 6.9 3.4 $\pm$ 1.4 71 $\pm$ 8.6
${}^{\dagger}$ [8] 91.4 $\pm$ 1.5 88.5 $\pm$ 1.4 86.5 $\pm$ 1.2 80.8 $\pm$ 1.9 2.5 $\pm$ 1.2 82 $\pm$ 6.9 3.4 $\pm$ 1.4 71 $\pm$ 8.6
$*{}^{*}$[6] 90.6 $\pm$ 0.9 93.9 $\pm$ 0.7 91.0 $\pm$ 2.0 93.7 $\pm$ 2.3
$*{}^{*}$[6] 90.4 $\pm$ 2.5 92.8 $\pm$ 1.0 91.9 $\pm$ 0.8 92.7 $\pm$ 1.2
@@
TABLE II: Comparison to State-of-the-Art on Synthetic data. Simulated results on the packed and piled scenes using the evaluation setup from [6]. We report grasp success rate (GSR), declutter rate (DR), and reconstruction performance, Chamfer L1 distance (C-L1), IoU. ICGNet denotes our approach that predicts grasp poses and occupancy for each individual instance. We retrain (${}^{\dagger}$) GIGA on randomized viewpoints and 2M grasps for each environment separately and re-evaluate ($*{}^{*}$) the pre-trained checkpoints provided by [6] multiple times. The scores for VGN and PointnetGPD are taken from [6]. Highest number marked in bold.
@@ IV Experiments
[…]
Figure 4: Our grasp predictions on simulated, unseen test objects from Google Scanned Objects[35]. Predicted grasps for “bottle”, “can” and “box” in the packed (left) and pile setup (right).
@@
Grasp Selection and Reconstruction.   For each scene, 32 instance queries are spawned and refined using our query refinement and the observed, denoised pointcloud. The pointcloud is downsampled using a voxel size of $2$ mm before computing the grasp affordances for each point. Surface normals for each contact point are estimated using covariance analysis from Open3D [37]. We then execute the grasp associated with the highest affordance score which is collision-free, given our implicit shape encodings of the scene and height of the table. To evaluate the scene reconstruction quality, we compute the occupancy field for the whole scene by combining instance level predictions $p^{scene}_{\mathtt{occ}}(x)=\max_{\mathcal{I}\in Inst}p_{\mathtt{occ}}(x|\mathcal{I})$ and use adaptive marching cubes algorithm [27].
@@
Simulation Results.   We evaluate our method using the evaluation process from [6] for the pile and packed scenes. For the grasping task, we report the mean and standard deviation of the GSR and DR over 4 different runs, each consisting of 100 different scenes. For the reconstruction task, we randomly sample 100 different pile and packed scenes using the objects from the test split. The Chamfer L1 and IoU scores are calculated following [27]. Scores are shown in Tab. II where ICGNet exhibits superior or equal performance compared to other methods in both grasp success and reconstruction, achieving the highest GSR, DR, IoU and lowest Chamfer L1 distance. Our method significantly outperforms the baseline methods on the packed dataset, highlighting the advantage of our grasp representation and instance priors. We additionally compare the model size and performance scores in I, showing that our model is competitive in size and inference speed while combining multiple tasks.
@@
Challenging Viewpoints.   We further evaluate our method on the challenging task of grasp detection from top-down viewpoints. To this end, we modify the evaluation algorithm from [6] to spawn the camera facing downwards at a random angle $\theta\in[0,2^{\circ}]$. This setup is extremely challenging for contact-based grasping since almost no graspable regions are observed and the surface normals are harder to estimate, resulting in significantly lower declutter rates of (85% and 72%) on the packed and pile environments respectively. To overcome this limitation of contact based grasping, we leverage our multi-task architecture and resample, unobserved points from the (implicit) object surfaces if no feasible grasp is found. We then use these new surface points for grasp prediction, allowing to improve the declutter rate to 90%(+6%) and 94% (+31%) respectively.
@@
Real World Experiments.   We validate our method and compare it against GIGA [8] and VN-EdgeGraspNet [6] on different real-world declutter tasks. The experimental setup is depicted in Fig. 5. A RealSense D415 RGB-D camera is attached to the gripper on a Franka Research 3 Arm. We use a total of 17 different test objects. Of these, 3-6 objects are randomly placed on a $30\times 30$ cm workspace. Before each experiment, we collect a top-down image of all objects, which are placed on a grid to ensure the repeatability and a fair comparison of all methods. Before each grasp trial, the arm moves to the designated image acquisition position (see Fig. 5) and captures depth image, which is post-processed and fed into the network. We use MoveIt! [38] to plan and execute collision-free grasp maneuvers. For collision checking, we rely on the predicted scene reconstructions or use the collision mesh from the tsdf volume. Given a set of predicted grasp poses, we filter the grasps for a 0.9 confidence level and execute the pose with the highest z component that is kinematically feasible and collision-free. If no grasp is found, we lower the confidence level to 0.8 and 0.7. The current task is terminated if two consecutive grasp failures occurs or if no feasible grasps are found for five sequential observations. Grasp success rate (GSR) and declutter rate (DR) for different clutter categories are reported in Tab. III.
@@
[…]
Limitations.   Although our architecture demonstrates remarkable performance, we have identified certain limitations. In line with [26], we observe that our network occasionally over-segments or combines instances even when they are not in contact. Most of these errors can be corrected by post-processing the network predictions, which comes at the cost of higher latency. Additionally, we observe that the reconstructions can be further improved by learning occupancy for the full scene instead of instances. We plan to address these limitations in future work by utilizing dilated attention to to improve spatial coverage of the query and include color into the pointcloud features. Additionally, we plan to train our network on datasets with diverse objects in order to learn robust object priors.

### v2

@@ I Introduction
In experiments, significant improvements arise from our object-centric architecture, establishing a new state-of-the-art in the packed decluttering benchmark introduced in [4] while surpassing scene-centric task baselines [8, 6]. These improvements are reflected in real-world experiments deployed on the Franka Research 3 robot. We additionally illustrate how instance-centric information can facilitate target-driven grasping (i.e., “grasp instance number 1" or “grasp the bottle") and effectively prevent post-grasp object-object collisions.
@@ II Related Work
Deep Grasp Synthesis.   Recent advancements in robotic grasping have predominantly utilized deep learning techniques to detect robotic grasps directly from sensor data. Unlike more traditional heuristics [16], these methods have demonstrated superior generalization performance towards previously unseen objects as well as achieving successful grasps in cluttered scenes. While several methods have achieved high success rates in 4-degree-of-freedom (4 DoF) top-down grasping [2, 17], their capabilities are often limited when dealing with cluttered scenes or task-dependent grasping scenarios [9]. This limitation has led to an increased research focus on 6 DoF grasp predictions, which aim to predict the full grasp pose from visual observations. We further classify these methods in sampling- and scene-based approaches. Sampling-based methods sample different grasp candidates and use a detector or diffusion [18, 19] network to refine them further. GPD [3] and PointNetGPD [20] learn to detect grasp poses in cluttered scenes from raw pointclouds by first sampling feasible grasp predictions and then using a CNN or PointNet classifier to predict the quality of each grasp based on the enclosed points. 6-DoF Graspnet [9] extends the way that grasp proposals are generated to the full SE(3) space by leveraging a variational auto-encoder for singulated objects. Scene-based approaches directly predict feasible grasp poses for the whole scene in one forward pass. Contact based methods such as Edge Grasp Networks [6] or Contact GraspNet [5] directly predict the SE(3) pose, width, and grasp quality for each point in the pointcloud. VGN [4] and GIGA [8] predict grasps for each voxel in the reconstructed TSDF using 3D CNN.
@@ III Method
Figure 2: Grasp Representation. Left: $\mathbf{c}$ is the contact point of the closed gripper and $\mathbf{n}$ is its estimated surface normal. $\mathbf{a}$ is the approach direction of the gripper. Given the gravity vector $\mathbf{z}$ and surface normal $\mathbf{n}$, $\mathbf{a}$ can be uniquely defined by the approach angle $\alpha$. Top Right: Grasp ambiguity of different grasp representations. When dealing with a particular contact or gripper center, there can be multiple feasible approach direction resulting in a successful grasp. Bottom Right: For each contact point, our representation enables the prediction of grasp qualities for different gripper orientations perpendicular to the surface normal.
@@
The corresponding set of SE(3) poses for a given grasp affordance prediction $g$ and gravity vector $\mathbf{z}$ is given by the tool-center point and computed as
@@
Encoder. Given a pointcloud in the world frame, our encoder network performs several key tasks. It extracts both sparse and dense features at multiple resolutions utilizing a sparse 3D-UNet architecture [26] for surface features and dense 3D-UNet architecture [27] for volumetric features. The sparse features are enriched with volumetric information through the proposed token aggregation module. To decompose the voxelized scene into individual instances, we apply multiple instance query refinement modules similar to Mask3D [28]. Further, we rely on a classification head to assign a class to each latent representation and filter out unmatched queries using a no-object class. All valid instance queries are directly used as latent embeddings for an occupancy network. Additionally, we employ a final Self Attention and MLP layer to exchange inter-object information between the queries and convert them to the affordance domain.
@@
Decoder. Our decoder is designed as an implicit neural field and uses world coordinates ($x,y,z$) as input. It predicts occupancy $p_{occ}$ and grasp affordances $g\in\mathcal{G}_{aff}$ for each instance. In addition, predictions such as classification and instance heatmaps are directly extracted from the encoder as shown in Fig. 3. These predictions are made for each instance query (class) and each point within the input pointcloud (instance id). To model each implicit decoder, we use a series of MLPs with residual connections following the approach proposed in [29]. Further, we concatenate each instance query with its positional encoding, which captures spatial information. Moreover, the queried coordinates ($x,y,z$) are enriched by incorporating the surface and volumetric features using our interpolation module. These enriched coordinates are then concatenated with the positional encodings of the original coordinates.
@@
Token Aggregation. Ideally, our goal is to extract dense features at a high resolution and directly feed them into the instance query refinement module. However, this approach comes with significant drawbacks, notably increased memory consumption and computation cost, rendering it infeasible for most applications. On the other hand, sparse neural networks scale with the number of occupied voxels, making them suitable for scaling to large areas. This advantage arises as the number of occupied voxels typically grows slower than the total number of voxels. Here, we combine the best of both by enriching each occupied surface feature with a volumetric context that is extracted from the volumetric feature grid. To this end, each surface feature is concatenated with the feature of the nearest volumetric voxel, similar to PointNet [30]. This enables the volumetric feature backbone to operate on a larger resolution extracting context information to enrich the sparse features used in the query refinement.
@@
Query Refinement. Given a set of $K$ instance queries, we apply a series of masked cross-attention and self-attention to extract instance-centric information given the scene-level features extracted from the U-Net backbones. We adapt Mask3D [28] and add an MLP layer to further process the extracted scene features. We also add Fourier positional encodings [31] based on voxel positions and use farthest-point sampling to sample initial instance query positions.
@@
Classifier. The classifier predicts a class label including a non-object class to address the varying number of instances in a scene. A small MLP followed by a softmax activation predicts a categorical distribution over the desired $C+1$ classes. For our experiments, we manually annotate the points with class labels that correlate with the respective shape, consisting of six categories (“mug", “box", “can", “bottle", “cylindric", “ball" and “other"). Having the notion of semantic classes, allows to predict grasp candidates and reconstructions for individual objects as shown in Fig. 4.
Loss Formulation.   The proposed model is trained in an end-to-end fashion. The training uses instance and semantic annotations as well as occupancy and grasp poses obtained from simulation. The implemented loss consists of binary cross-entropy, DICE and squared-error losses and is given as
@@
where the first three terms refer to the panoptic segmentation task and are calculated with respect to the input pointcloud $P_{\mathtt{scan}}$. $\mathcal{L}^{Grasp}$ refers to the grasp predictions for each instance. This loss is calculated with respect to the sampled grasp coordinates which may differ from the actual observed pointcloud. We use the cross entropy loss to supervise the approach classification and the squared L2 loss for the gripper width. Finally, the reconstruction loss $\mathcal{L}_{\mathtt{occ}}^{BCE}$ supervises the predicted occupancy for each instance and is computed with respect to an additional set of randomly sampled points within the scene. More information about the labels and dataset is provided in Sec. IV. Note that there is no ordering in the set of instances in a scene and we need to establish correspondence between the set of predicted and set of groundtruth instances during training. Following [28], we use Hungarian matching based on the instance segmentation [^3] We limit the matching to the segmentation task since calculating every possible assignment of the occupancy field is computationally extensive. loss to find unique assignments and apply the same loss function to the predictions at multiple resolutions.
@@
GIGA [8] 89.9 $\pm$ 1.7 87.6 $\pm$ 2.0 76.3 $\pm$ 2.4 80.9 $\pm$ 4.1 2.5 $\pm$ 1.2 82 $\pm$ 6.9 3.4 $\pm$ 1.4 71 $\pm$ 8.6
GIGA-HR [8] 91.4 $\pm$ 1.5 88.5 $\pm$ 1.4 86.5 $\pm$ 1.2 80.8 $\pm$ 1.9 2.5 $\pm$ 1.2 82 $\pm$ 6.9 3.4 $\pm$ 1.4 71 $\pm$ 8.6
EdgeGraspNet[6] 90.6 $\pm$ 0.9 93.9 $\pm$ 0.7 91.0 $\pm$ 2.0 93.7 $\pm$ 2.3
VN-EdgeGraspNet[6] 90.4 $\pm$ 2.5 92.8 $\pm$ 1.0 91.9 $\pm$ 0.8 92.7 $\pm$ 1.2
@@
TABLE II: Comparison to State-of-the-Art on Synthetic data. Simulated results on the packed and piled scenes using the evaluation setup from [6]. We report grasp success rate (GSR), declutter rate (DR), and reconstruction performance, Chamfer L1 distance (C-L1), IoU. ICGNet denotes our approach that predicts grasp poses and occupancy for each individual instance. We retrain () GIGA on randomized viewpoints and 2M grasps for each environment separately and re-evaluate () the pre-trained checkpoints provided by [6] multiple times. The scores for VGN and PointnetGPD are taken from [6]. Highest number marked in bold, second highest underlined.
@@ IV Experiments
Simulation Environment and Dataset.   We leverage the simulation setup and the object dataset introduced in VGN [4]. This dataset consists of 303 train and 40 test objects from various sources [34, 35, 36]. The experimental setup involves a free-floating Franka Emika Gripper and we sample grasps and occupancy values in $30$ cm3 sized tabletop workspace. Contrary to  [4, 8], which randomly samples the gripper centers close to the surface and stores the resulting best gripper orientation as a ground truth label, we adapt a different strategy. We select an observed point and corresponding surface normal, then execute the grasp from different approach angles as outlined in Sec. 3. Specifically, we sample twelve different approach angles for each contact point and the final gripper width as well as the grasped object are recorded. Additionally, we sample the occupancy values for each scene. A total of 200’000 occupancy values are sampled of which 70% are uniformly sampled and 30% are sampled closer to the object surfaces to allow more accurate reconstruction. We store the occupancy value for each object yielding a total of 200’000 $\times$ $k$ binary labels for each scene. We further make use of the packed and piled splits introduced by [4] and sample a total of 1M and 2M grasps from each split, resulting in 3M grasps from 15’000 scenes of which 13’500 and 1’500 are used for training and validation respectively.
@@
Grasp Selection and Reconstruction.   For each scene, 32 instance queries are spawned and refined using our query refinement and the observed pointcloud. The pointcloud is denoised using statistical outlier filters. It is then downsampled using a voxel size of $2$ mm before computing the grasp affordances for each point. Surface normals for each contact point are estimated using covariance analysis from Open3D [38]. We then execute the grasp associated with the highest affordance score which is collision-free, given our implicit shape encodings of the scene and height of the table. To evaluate the scene reconstruction quality, we compute the occupancy field for the whole scene by combining instance level predictions $p^{scene}_{\mathtt{occ}}(x)=\max_{\mathcal{I}\in Inst}p_{\mathtt{occ}}(x|\mathcal{I})$ and use adaptive marching cubes algorithm [29].
Figure 4: Our grasp predictions on simulated, unseen test objects from[39]. Predicted grasps for “bottle”, “can” and “box” in the packed (left) and pile setup (right). More qualitative examples at can be found on the project page
@@
Simulation Results.   We evaluate our method using the evaluation process from [6] for the pile and packed scenes. For the grasping task, we report the mean and standard deviation of over 4 different runs, each consisting of 100 different scenes. We report the Grasp success rate (GSR, percentage of successful grasps) and declutter rate (DR, percentage of removed objects after the task has finished) For the reconstruction task, we randomly sample 100 different pile and packed scenes using the objects from the test split. The Chamfer L1 and IoU scores are calculated following [29]. Scores are shown in Tab. II where ICGNet exhibits superior or equal performance compared to other methods in both grasp success and reconstruction, achieving the highest GSR, DR, IoU and lowest Chamfer L1 distance. Our method significantly outperforms the baseline methods on the packed dataset, highlighting the advantage of our grasp representation and instance priors. We additionally compare the model size and performance scores in I, showing that our model is competitive in size and inference speed while combining multiple tasks.
@@
Challenging Viewpoints.   We further evaluate our method on the challenging task of grasp detection from top-down viewpoints. To this end, we modify the evaluation algorithm from [6] to spawn the camera facing downwards at a random angle $\theta\in[0,2^{\circ}]$. This setup is extremely challenging for contact-based grasping since almost no graspable regions are observed and the surface normals are harder to estimate, resulting in significantly lower declutter rates of (85% and 72%) on the packed and pile environments respectively. To overcome this limitation of contact based grasping, we leverage our multi-task architecture and resample, unobserved points from the (implicit) object surfaces if no feasible grasp is found. We then use these new surface points for grasp prediction, allowing us to improve the declutter rate to 90%(+6%) and 94% (+31%) respectively.
@@
Real World Experiments.   We validate our method and compare it against GIGA [8] and VN-EdgeGraspNet [6] on different real-world declutter tasks. The experimental setup is depicted in Fig. 5. A RealSense D415 RGB-D camera is attached to the gripper on a Franka Research 3 Arm. We use a $30\times 30$ cm workspace to allow for comparison with  [8]. [^4] Note that our method is capable of inferring grasps on larger scenes due to the combination of dense and sparse voxelization. We use a total of 17 different test objects. Of these, 3-6 objects are randomly placed on a $30\times 30$ cm workspace. Before each experiment, we collect a top-down image of all objects, which are placed on a grid to ensure the repeatability and a fair comparison of all methods. Before each grasp trial, the arm moves to the designated image acquisition position (see Fig. 5) and captures depth image, which is post-processed and fed into the network. We use MoveIt! [40] to plan and execute collision-free grasp maneuvers. For collision checking, we rely on the predicted scene reconstructions or use the collision mesh from the tsdf volume.Given a set of predicted grasp poses, we follow the procedure of [6] and filter the grasps for a 0.9 confidence level and execute the pose with the highest z component that is kinematically feasible and collision-free. If no grasp is found, we lower the confidence level to 0.8 and 0.7. The current task is terminated if two consecutive grasp failures occur or if no grasps are found for five sequential observations. Grasp success rate (GSR) and declutter rate (DR) for different clutter categories are reported in Tab. III.
@@
Figure 5: Real world experimental setup. We use 17 different objects of which 3-6 are placed on a $30$ cm3 workspace.
Limitations.   Although our architecture demonstrates remarkable performance, we have identified certain limitations. In line with [28], we observe that our network occasionally over-segments or combines instances even when they are not in contact. Most of these errors can be corrected by post-processing the network predictions, which comes at the cost of higher latency. Additionally, we observe that the reconstructions can be further improved by learning occupancy for the full scene instead of for each instance independently. We plan to address these limitations in future work by utilizing dilated attention to to improve spatial coverage of the query and include color into the pointcloud features. Additionally, we plan to train our network on datasets with diverse objects in order to learn robust object priors.